OpenStudy (anonymous):

Why \[ \lim_{n \to \infty}|\frac{n^{2}+2n+1}{n^{2}}|=1\] ?

hartnn (hartnn):

1+2/n+1/n^2 if u put n=infinity 2/n=0 1/n^2=0 thats why.

OpenStudy (anonymous):

Because when you work out the division you have: \[1+\frac{2}{n}+\frac{1}{n^2} \] As n tends to infinity, the second 2 terms tend to 0, and we are left with 1.

OpenStudy (anonymous):

thanks i get it guys

OpenStudy (anonymous):

n^2 + 2n+1= (n+1)^2 so (n+1)^2/n^2 (n